Supposed that ln 2=a and ln 5=b. Use the properties of logarithms to write each logarithm of a and b. ln 32
I'm not sure where to begin this problem.

Answer by stanbon(75887)

Supposed that ln 2=a and ln 5=b. Use the properties of logarithms to write each logarithm of a and b. ln 32
----------------
ln(32) = ln(2^5) = 5*ln(2) = 5a
=====================================
Cheers,
Stan H.